Some chemical reactants are listed in the table below. Complete the table by filling in the n state of the highlighted atom. oxidation state of highlighted atom CO3 (aa) HPO () NOs) O Type here to search
Answer #1

CO3^2-

carbon oxidation state let x

x+ -2*3 = -2

x = 4

The oxidation state of carbon in CO3^2- is +4

HPO4^2-(aq)

the oxidation state of phosphorous is let x

1+x+(-2*4) = -2

   x = 5

the oxidation state of phosphorous is +5

VO(s)

The oxidation sate of V is let x

x-2 = 0

x   = +2

The oxidation state of V in VO is +2

N2(g)

free sate molecules oxidation sate is zero

The oxidation state of nitrogen in N2 is 0